# Deep Learning Accelerated Quantum Transport Simulations in Nanoelectronics: From Break Junctions to Field-Effect Transistors

### v1

@@ III.2 Double-gated Carbon Nanotube Field-effect Transistor
Having demonstrated the capability of DeePTB-NEGF in simulating quantum transport in break junction systems, we now extend our framework to semiconductor devices where the interplay between quantum transport and electrostatic effects becomes crucial. Here we focus on carbon nanotube field-effect transistors (CNT-FETs), which are promising candidates for sub-5 nm technology nodes due to their excellent carrier transport properties[30, 31]. The quasi-one-dimensional structure and long mean-free path of CNT enable near-ballistic transport at nanoscale dimensions, making them ideal channel materials for high-performance electronic devices[32]. The realistic simulation of CNT-FETs requires addressing the electrostatic effects on quantum transport under finite gate voltage and drain-source bias, which can be achieved through the NEGF-Poisson approach[33, 34].
While DFT-NEGF methods have been successful in quantum transport calculations, they become computationally prohibitive for experimentally relevant device dimensions. The TB method offers a more efficient alternative by employing smaller and sparser Hamiltonian matrices. Here we implement the DeePTB model which enhances traditional TB approaches by incorporating local environment-dependent corrections to the SK parameters (Eq.1). This approach goes beyond the conventional two-center approximation while maintaining the computational efficiency of TB methods[12]. Combined with the NEGF-Poisson formalism, our method enables efficient self-consistent quantum transport calculations for realistic devices under finite bias conditions.
The self-consistent calculation begins with the equilibrium TB Hamiltonian $H_{0}$ and iteratively solves for the non-equilibrium potential profile using the Poisson equation [15, 35]:
$\nabla[\epsilon(\vec{r})\nabla\phi(\vec{r})]=-e[\rho_{\text{free}}(\vec{r})+\rho_{\text{fix}}(\vec{r})]$ (6)
@@
Here, $\epsilon$ is the dielectric constant, $\phi$ is the electrostatic potential, $\rho_{\text{free}}$ is the free charge concentration, and $\rho_{\text{fix}}$ represents the fixed charge from doping. Specifically, $\rho_{\text{fix}}$ corresponds to the average density of donated or accepted electrons per atomic site in the doped region. The NEGF-Poisson self-consistency is achieved through the Gummel iteration scheme [36], with the Poisson equation solved using the Newton-Raphson method on a discrete real-space grid. Within our TB-NEGF framework, the spatial charge distribution is approximated by Mulliken charges localized at atomic sites.
To demonstrate the capability of our implementation, we study quantum transport in a double-gated CNT-FET, as illustrated in Fig. 4(a). The device consists of a carbon nanotube channel with p-doped source and drain regions ($\rho_{\text{fix}}=5\times 10^{-3}$ per atomic site), sandwiched between top and bottom gates. The gates encompass the entire channel length $L_{\text{G}}$ and are separated from the CNT by dielectric layers of thickness $t_{\text{ox}}=1.7$ nm. The electronic structure is described using the DeePTB model implemented with the conventional Slater-Koster parameters ($h_{ll^{\prime}{\zeta}}(r_{ij})$ term in Eq. 1) without environmental corrections, to enable direct comparison with established TB-NEGF codes. A simplified $\pi$-bond model is adopted with one $p_{z}$ orbital per carbon atom and hopping energy $t=1.09$ eV. The electrostatic boundary conditions are set as follows: Dirichlet conditions at the gate regions (fixed to gate voltage), von Neumann conditions (zero electric field) at the exposed dielectric surfaces, and flat potential profiles with von Neumann conditions at the edges of the semi-infinite source and drain regions. The dielectric environment is controlled through the spatial distribution of the dielectric constant $\epsilon(\vec{r})$.
We first validate our NEGF-Poisson implementation by comparison with the established TB-NEGF code NanoTCAD ViDES[15, 16]. Under identical simulation conditions ($L_{\text{G}}=2$ nm, $V_{\text{gs}}=1$ V, and $V_{\text{ds}}=0.5$ V), both the transmission spectra (Fig.4(b)) and electrostatic potential energy distribution (Fig.4(c)) show excellent agreement. The difference in potential energy distribution (Fig.4(d)) shows a maximum deviation of less than $7\times 10^{-3}$ eV, confirming the accuracy of our self-consistent implementation in the DeePTB-NEGF framework.
Having validated our implementation, we investigate the device transfer characteristics under different operating conditions, focusing on the effects of dielectric materials and gate lengths. Fig. 4(e) compares the transfer characteristics with two different gate dielectric materials, SiO2 and HfO2, at $V_{\text{ds}}=0.5$ V and $L_{\text{G}}=2$ nm. For direct comparison, the gate work functions are adjusted to achieve the same off-current $I_{\text{OFF}}=2.5\times 10^{-7}$ A at $V_{\text{gs}}=0.45$ V. When switching from SiO2 to HfO2(dielectric constant changing from $3.9$ to $20$ in gate dielectric area), the subthreshold swing (SS), which characterizes the inverse slope of the transfer characteristics curve on a semi-logarithmic scale, decreases from 696.56 $696.56\text{ mV dec}^{-1}$ to 213.72 $213.72\text{ mV dec}^{-1}$. This improvement aligns with the fact that high-$\kappa$ gate dielectrics reduce the effective screening length of the channel $\lambda_{\text{ch}}\propto\sqrt{(\epsilon_{\text{ch}}/\epsilon_{\text{ox}})t_{\text{ox}}t_{\text{ch}}}$ where $\epsilon_{\text{ch}}$ and $\epsilon_{\text{ox}}$ are the dielectric constants for channel and gate dielectric, and $t_{\text{ch}}$ and $t_{\text{ox}}$ represent their respective thicknesses[37]. The higher $\epsilon_{\text{ox}}$ of HfO2 leads to shorter effective screening length, indicating stronger gate electrostatic control over the channel. The effect of gate length is investigated using SiO2 as the dielectric material (Fig. 4(f)). Increasing the gate length from 2 nm to 10.5 nm improves the SS from $449.47$ mV dec-1 to $92.12$ mV dec-1. This significant improvement occurs because longer gates provide better electrostatic control with larger gate capacitance.
The excellent agreement between our implementation and NanoTCAD ViDES validates our NEGF-Poisson self-consistent framework. The physically reasonable response to different dielectric materials and gate lengths further demonstrates its capability to capture essential device physics. These results establish DeePTB-NEGF as a versatile approach for quantum transport simulations in realistic semiconductor devices, combining the efficiency of TB methods with the accuracy needed for device design and optimization under finite bias conditions.
@@ IV Summary and Discussion
We have developed a deep learning-accelerated framework, DeePTB-NEGF, for efficient quantum transport calculations in nanoelectronic devices. By integrating deep learning Hamiltonian prediction with the NEGF method, our approach enables both direct quantum transport calculations and self-consistent solutions incorporating electrostatic effects. Most importantly, this framework achieves comprehensive simulation of the dynamic break junction process, extending to realistic semiconductor device simulations through NEGF-Poisson self-consistency. This versatility, combined with the method’s efficiency and accuracy, bridges the gap between atomic-scale quantum transport simulations to statistical analysis of experimental measurements and practical device applications.
The broad applicability and accuracy of this framework have been rigorously validated through two representative applications. In break junction systems, we demonstrate the capability of our method to comprehensively simulate the junction evolution in both the metallic contacts and single-molecule junction systems. Our predicted conductance histograms show good agreement with experimental measurements in both cases, reproducing characteristic features such as the quantized conductance peaks between $1.0\mathrm{G}_{0}$ and $2.0\mathrm{G}_{0}$ for metallic contacts and the pronounced peak at $10^{-3.3}\mathrm{G}_{0}$ for molecular junctions. Building upon this success in break junction systems, the framework’s extension to semiconductor devices is validated through carbon nanotube FET simulations, where our NEGF-Poisson implementation accurately captures device electrostatics and transport characteristics under various conditions, as confirmed by comparison with established quantum transport codes.
Based on these achievements, the demonstrated performance of DeePTB-NEGF establishes a new paradigm for high-throughput quantum transport calculations in nanoelectronics. The framework’s dual capability in simulating break junction processes and transistors offers opportunities to advance both molecular electronics research and practical device engineering. Further development of this methodology could incorporate more sophisticated quantum effects and extend to larger-scale systems, potentially enabling quantum transport simulations at industrially relevant scales. The successful integration of deep learning efficiency with quantum mechanical accuracy presents a significant step toward bridging fundamental physics and technological applications in nanoelectronics.
Code availability
The DeePTB-NEGF code is openly available in the github deepmodeling community at https://github.com/deepmodeling/DeePTB. For data preparation, we also developed and released an open-source tool dftio.
Acknowledgements.

### v2

@@ III.2 Carbon Nanotube Field-effect Transistor
Having demonstrated the capability of DeePTB-NEGF in simulating quantum transport in break junction systems, we now extend this framework to field-effect transistors (FETs). The critical distinction in FETs is the presence of gate-modulated electrostatics, which necessitates a self-consistent solution of the Poisson equation coupled with the NEGF formalism [29, 30]. This extension enables the simulation of more complex nanoelectronic devices where the interplay between quantum transport and electrostatic effects is paramount. We apply this enhanced methodology to carbon nanotube field-effect transistors (CNT-FETs), which are promising candidates for next-generation electronics due to their exceptional charge transport properties [31, 32, 33]. The quasi-one-dimensional structure and long mean free path of CNTs facilitate near-ballistic transport at nanoscale dimensions, making them ideal for high-performance devices [34]. Accurate modeling of CNT-FETs requires capturing the gate-voltage-dependent modulation of carrier transport, a challenge well-suited to our extended DeePTB-NEGF approach.
To incorporate gate electrostatics in CNT-FET simulations, we implement the NEGF-Poisson formalism using TB models predicted by DeePTB, which provides local environment-dependent corrections to the Slater-Koster parameters (Eq. 1), offering advantages over DFT Hamiltonian models due to their smaller, sparser matrices and direct compatibility with the NEGF-Poisson framework. This self-consistency is achieved through the Gummel iteration scheme [35], which begins with the equilibrium TB Hamiltonian $H_{0}$ and iteratively solves for the non-equilibrium potential profile using the Poisson equation [36, 37]
$\nabla[\epsilon(\vec{r})\nabla\phi_{\text{n+1}}(\vec{r})]=-e[\rho_{\text{n}}(\vec{r})+\rho^{\text{fix}}(\vec{r})].$ (6)
@@
with the Newton-Raphson method on a discrete real-space grid. Here $\epsilon$ is the dielectric constant, $\phi_{\text{n+1}}$ and $\rho_{\text{n}}$ are the electrostatic potential and free charge number concentration in the $n$-th iteration. The term $\rho^{\text{fix}}$ represents the average equivalent charge associated with ionized donors or acceptors at each atomic site within the doped region. In practical calculations, $\rho_{\text{n}}$ is obtained as Mulliken charge at all atomic sites based on the Hamiltonian $H_{\text{n}}$ from the n-th iteration . The boundary conditions for the Poisson equation include constraints from gate voltage $V_{\text{gs}}$ and drain-source bias $V_{\text{ds}}$. Within DeePTB-NEGF framework, the predicted TB Hamiltonian is updated according to the electrostatic potential as:
$H_{n}^{i}=H_{0}^{i}-e\phi_{\text{n}}({\bm{r}}_{i})$ (7)
where $H_{n}^{i}$ and $\phi_{\text{n}}({\bm{r}}_{i})$ are the $n$-th iteration Hamiltonian onsite block and electrostatic potential at site $i$, respectively. The complete NEGF-Poisson procedure is detailed in Sec.S2.B in SM. Our implementation has been validated against the established TB-NEGF code NanoTCAD ViDES [36, 38], as described in SM Sec.S2.C.
We now apply the DeePTB-NEGF framework to study the transport properties of a local bottom gate (LBG) CNT-FET with experimental dimensions, demonstrating the capability of our method to simulate realistic nanoelectronic devices that are challenging for traditional first-principles approaches due to their large size. The LBG CNT-FET geometry, shown in Fig. 4(a), has been experimentally realized with sub-10 nm channel lengths [15]. This device exhibits remarkable performance, including an ON-OFF current ratio ($I_{\text{on}}/I_{\text{off}}$) of $10^{4}$ and a subthreshold slope (SS) of 94 mV/decade for a 9 nm channel length with Pd electrodes. While numerous theoretical studies have explored the transport properties of CNT-FETs using NEGF, most are based on parameterized Hamiltonians [39, 40, 41, 15] or simulated with thinner tubes than those used in experiments due to computational limitations [42]. These approaches lack sufficient accuracy in electronic structure or direct comparability with experimental results at the first-principles level without size approximations. By simulating this specific device with DeePTB-NEGF framework, we aim to validate our method against experimental results and explore the scaling limits of CNT-FETs. This approach enables us to address challenges that are difficult for both experiments and conventional simulation techniques while avoiding inaccuracies introduced by smaller geometries when solving Poisson equation and determining charge density.
To investigate the influence of CNT diameter on device performance, we simulated CNT-FETs with various indices $(10,0),(13,0)$, $(16,0)$ and $(17,0)$ (diameters ranging from 0.79 nm to 1.35 nm). This range encompasses the  1.3 nm diameter CNTs used in experimental devices [15], with CNT(16,0) and (17,0) (diameters 1.27 nm and 1.35 nm, respectively) being most comparable to the experimental sample. We trained a $sp$-TB model with environment-dependent SK parameter using the DFT eigenvalues of the CNT$(7,0),(10,0),(13,0)$ and $(16,0)$ structures as training data set, and then the band of CNT$(17,0)$ can be predicted. The details about data preparation and model training are provided in the SM Sec.S2.A. A comparison of band structures for these CNTs between DFT and DeePTB calculations are illustrated in the Fig. S2 in SM, validating the accuracy and generality of DeePTB model across CNTs with different diameters. To model the metal contacts, we employed a doping contact approach by setting $\rho_{\text{fix}}(r)$ in Poisson equation (Eq. 6). This method, used in previous theoretical studies [15, 42], induces a built-in electrostatic field between the doped regions and channel mimicking the band bending at the Pd electrode-CNT interface in Ohmic contacted CNT-FET [43, 44]. We tested doping concentrations ranging from $5.33\times 10^{8}\text{m}^{-1}$ to $1.25\times 10^{9}\text{m}^{-1}$, with results showing high consistency in ON current and SS (see Sec. S2.D in SM). Based on these tests, we chose a doping concentration of $8.89\times 10^{8}\text{m}^{-1}$ for subsequent simulations. Additionally, the bottom gate covers the channel as well as the sources and drain contacts in LBG device. Therefore, we set the total gate length as $L_{\text{tot}}=L_{\text{ch}}+2\times L_{\text{ext}}$ , including the channel length $L_{\text{ch}}$ and source/drain extensions $L_{\text{ext}}$[15]. In order to ensure the bulk-like environment inside the electrodes, different $L_{\text{ext}}$ have been tested in Sec. S2.D in SM. It indicates that $L_{\text{ext}}=2.5\text{nm}$ is long enough to screen the impact from channel. The other detailed simulation settings are listed in Sec. S2.D in SM.
Figure 4(b) shows the simulated transfer characteristics for CNTs with different diameters, alongside experimental data. For clear comparison, the gate work functions of all devices have been shifted to have the same ON current $I_{\text{on}}\approx 1\times 10^{-5}$ A at $V_{\text{gs}}=-1.0$ V. The numbers of atoms in these devices with $9$ nm channel length are 1620, 2106, 2592 and 2754 for CNT (10,0), (13,0), (16,0) and (17,0) respectively, which are beyond the computing ability of traditional first-principles NEGF-Poisson SCF. As illustrated in Fig. 4(b), the transfer characteristics for the four tubes are similar in the ON state while becoming drastically different when entering the subthreshold region. These differences arise from the distinct band gaps of CNTs with varying diameters. For the same p-doped tube, the barrier height between the source and channel increases as the gate voltage $V_{\text{gs}}$ becomes more positive, hindering the injection of the holes from the source. This behavior is illustrated by the local density of states (LDOS) of CNT$(17,0)$ shown in Fig. 4(d). Despite the differences in diameter, $I_{\text{on}}$ remains stable at approximately $1\times 10^{-5}$ A, attributed to similar transmission functions inside the drain-source bias window in the ON state without barrier hindering (see Sec.S2.D in SM). However, the OFF current $I_{\text{off}}$ decreases with diameter shrinkage because a larger band gap in thinner CNT tends to induce a higher potential barrier $\Phi_{b}$ at OFF state. The barrier suppresses source-drain tunneling, which is the major contribution to $I_{\text{off}}$. [40, 45]. LDOS calculations can reveal a significant decrease in $\Phi_{b}$ with increasing CNT diameter(see Fig. S6 in SM). At OFF states, $\Phi_{b}$ reduces from 0.55 eV for CNT(10,0) to 0.19 eV for CNT(17,0). Here the OFF state for CNT(10,0) is taken as $V_{\text{gs}}=0.3$V with $I_{\text{ds}}<10^{-13}$. This trend is consistent with the observed increases in $I_{\text{off}}$. This comprehensive study of CNT diameter effects demonstrates the importance of accurately matching the simulated CNT size to experimental conditions for meaningful transfer characteristics and figures of merits such as $I_{\text{on}}/I_{\text{off}}$.
To investigate the scaling limit of LBG CNT-FET, we simulate the transfer characteristics using CNT(16,0) for device with various channel lengths. The scalability of FETs can be quantified by the natural length $\lambda=\sqrt{(\epsilon_{\text{ch}}/n\epsilon_{\text{ox}})t_{\text{ox}}t_{\text{ch}}}$ where $n$ is the equivalent number of gates, $\epsilon_{\text{ch}}$ and $\epsilon_{\text{ox}}$ are the dielectric constants for channel and gate dielectric, and $t_{\text{ch}}$ and $t_{\text{ox}}$ represent their respective thicknesses [46]. For good electrostatics to avoid short channel effect(SCE), channel length $L_{\text{ch}}>(3\sim 5)\lambda$ is required [33]. In LBG geometry, $\lambda\approx 0.956$nm(detialed derivation in SM Sec.S2.D). Hence we explore the geometry with channel length $L_{\text{ch}}$ ranging from $9$ nm to $4$ nm, where the extent of short-channel effects (SCE) remains uncertain. As illustrated in Fig. 4(c), when $L_{\text{ch}}$ becomes shorter the SS increases from 98.88 mV/decade to 257.27 mV/decade with $I_{\text{on}}/I_{\text{off}}$ decreasing from $1\times 10^{5}$ to $4\times 10^{3}$. Obviously, it is a manifestation of SCE which can be explained by the gate electrostatic efficiency degradation in FET [33]. From the perspective of capacitance model, the gate electrostatic efficiency $\Delta V_{\text{g}}/\Delta\phi_{\text{ch}}$, the ability for gate voltage $V_{\text{g}}$ to tune the channel potential $\phi_{\text{ch}}$, can be expressed as
$\frac{\Delta V_{\text{g}}}{\Delta\phi_{\text{ch}}}=1+\frac{C_{\text{s}}+C_{\text{d}}+C_{\text{dep}}+C_{\text{it}}+C_{\text{Q}}}{C_{\text{gox}^{\prime}}}.$ (8)
Here $C_{\text{gox}^{\prime}}$ is the effective gate capacitance, $C_{\text{s}}$ and $C_{\text{d}}$ are parasitic capacitances from source and drain, $C_{\text{dep}},C_{\text{it}},C_{\text{Q}}$ are capacitances induced by substrate, interface traps and mobile charge carriers respectively. With shorter channel $C_{s}$ and $C_{d}$ are becoming comparable to $C_{\text{gox}^{\prime}}$, a result of the gate sharing control over the channel potential $\phi_{\text{ch}}$ with the source and drain in short-channel devices.[33]. Larger gate potential difference $\Delta V_{\text{g}}$ is needed to tune the same change in $\phi_{\text{ch}}$, thus degrading SS. The enhancing tendency of short channel effect is also reflected by LDOS (See Fig. S7 in SM). Under identical gate and drain-source bias conditions, a shorter channel device exhibits a lower and narrower barrier, leading to substantial source-drain tunneling and an increase in $I_{\text{off}}$ and SS [40]. Therefore, the scaling potential in the LBG geometry with a 3 nm $\text{HfO}_{2}$ dielectric thickness (used in our simulation and the experiment [15]) is limited, necessitating further improvements in device design. Additionally, the transfer characteristics of LBG CNT-FETs with varying channel lengths, shown in Fig. 4(c), emphasize the critical importance of simulated device dimensions in accurately reproducing the experimentally measured values, as previously discussed.
These results demonstrate the high fidelity of the DeePTB-NEGF framework in simulating quantum transport in CNT-FETs, achieving DFT-level accuracy with TB efficiency. The framework accurately captures device physics across various nanotube diameters and channel lengths, showing excellent agreement with experimental data. This approach enables quantum transport analysis in realistic semiconductor devices and facilitates the optimization of future nanoelectronic device designs.
@@ IV Summary and Discussion
In this work, we have developed DeePTB-NEGF, a deep learning-accelerated framework for efficient quantum transport simulations in nanoelectronic devices. This approach integrates deep learning (TB) Hamiltonian prediction with the NEGF method, enabling both direct quantum transport calculations and self-consistent solutions incorporating electrostatic effects. The framework achieves comprehensive simulation of dynamic break junction processes and CNT-FETs of experimental size under finite bias through NEGF-Poisson self-consistency.
The broad applicability and accuracy of this framework have been rigorously validated through the two representative applications. In break junction systems, we demonstrate the capability of our method to comprehensively simulate the junction evolution in both the metallic contacts and single-molecule junction. Our predicted conductance histograms show good agreement with experimental measurements in both cases, reproducing characteristic features such as the quantized conductance peaks between $1.0\mathrm{G}_{0}$ and $2.0\mathrm{G}_{0}$ for metallic contacts and the pronounced peak at $10^{-3.3}\mathrm{G}_{0}$ for molecular junctions. Extending the framework to semiconductor devices, we simulated transfer characteristics of LBG CNT-FETs with dimensions matching experimental devices using the DeePTB-NEGF-Poisson implementation. Our simulations achieved excellent agreement with the measured transfer characteristic curve, highlighting both the efficiency and accuracy of our method and underscoring the importance of device-level simulations. Furthermore, the scaling potential of the LBG device geometry is explored with channel lengths ranging from 9 nm to 4 nm. Our simulations demonstrate that, without improvements in the geometry, the scaling potential remains limited.
Based on these achievements, the demonstrated performance of DeePTB-NEGF establishes a new paradigm for high-throughput quantum transport calculations in nanoelectronics. The framework’s dual capability in simulating break junction processes and transistors in experimental size offers opportunities to advance both molecular electronics research and semiconductor device engineering. Further development of this methodology could incorporate more sophisticated quantum effects and extend to larger-scale systems, potentially enabling quantum transport simulations at industrially relevant scales. The successful integration of deep learning efficiency with quantum mechanical accuracy presents a significant step toward bridging fundamental physics and technological applications in nanoelectronics.
V Data availability
The DeePTB-NEGF code is openly accessible through the DeepModeling community’s GitHub repository at https://github.com/deepmodeling/DeePTB. All data supporting the findings of this study are available from the corresponding author upon reasonable request.
VI Acknowledgments